# Preferential Attachment Trees with Vertex Death: Persistence of the Maximum Degree

### v1

@@ 4.3. Preliminary results
We first collect a few preliminary results that are used throughout the proof of Theorems 4.2, 4.4, and 4.5.
@@
$\mathbb{P}\bigg{(}\exists j\in\mathbb{N}_{0}:Y+\sum_{i=1}^{k+j}X(i)\leq\sum_{i=k}^{k+j}X(i)^{\prime}\bigg{)}\leq\bigg{(}\prod_{i=k}^{\infty}\mathbb{E}\Big{[}\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\Big{]}\bigg{)}\mathbb{E}\bigg{[}\exp\bigg{(}-\lambda\bigg{(}Y+\sum_{i=1}^{k-1}X(i)\bigg{)}\bigg{)}\bigg{]}.$ (4.28)
Lemma 4.9 (Theorem $1.13$ in [24]).
Suppose that $(S_{i})_{i\in\mathbb{N}}$ is a sequence of mutually independent random variables such that $\sum_{i=1}^{\infty}S_{i}$ converges almost surely. The distribution of $\sum_{i=1}^{\infty}S_{i}$ contains an atom on $\mathbb{R}$ if and only if for some collection $(c_{i})_{i\in\mathbb{N}}\in\mathbb{R}^{\infty}$,
$\forall i\in\mathbb{N}:\mathbb{P}\!\left(S_{i}=c_{i}\right)>0\qquad\text{and}\qquad\sum_{i=1}^{\infty}\mathbb{P}\!\left(S_{i}\neq c_{i}\right)<\infty.$ (4.29)
@@
In case the $S_{i}$ are symmetric random variables, i.e. $\mathbb{P}\!\left(S_{i}\geq x\right)=\mathbb{P}\!\left(S_{i}\leq-x\right)$ for all $x\geq 0$ and $i\in\mathbb{N}$, the statement can be simplified by restricting to collections $(c_{i})_{i\in\mathbb{N}}$ that have finitely many non-zero values only. Indeed, let $(i_{j})_{j\in\mathbb{N}}$ be a strictly increasing sub-sequence such that $c_{i_{j}}\neq 0$ for all $j\in\mathbb{N}$. The first condition in (4.29) yields $\mathbb{P}\!\left(S_{i_{j}}=-c_{i_{j}}\right)=\mathbb{P}\!\left(S_{i_{j}}=c_{i_{j}}\right)>0$. Then, the condition
$\sum_{i=1}^{\infty}\mathbb{P}\!\left(S_{i}\neq c_{i}\right)<\infty$ (4.30)
in (4.29) implies that $\mathbb{P}\!\left(S_{i_{j}}=c_{i_{j}}\right)=1-o(1)$, where the $o(1)$ tends to zero with $j$. There thus exists $J\in\mathbb{N}$ such that $\mathbb{P}\!\left(S_{i_{j}}=c_{i_{j}}\right)=\mathbb{P}\!\left(S_{i_{j}}=c_{i_{j}}\right)>\frac{1}{2}$ for all $j\geq J$, which yields a contradiction. So, any collection $(c_{i})_{i\in\mathbb{N}}$ can only contain finitely many non-zero values.
Lemma 4.11 (Proposition $1.15$ in [24]).
Suppose that $(S_{i})_{j\in\mathbb{N}}$ is a sequence of mutually independent, symmetric random variables. If $\sum_{i=1}^{\infty}S_{i}$ diverges almost surely,
$\limsup_{n\to\infty}\sum_{i=1}^{n}S_{i}=\infty\qquad\text{and}\qquad\liminf_{n\to\infty}\sum_{i=1}^{n}S_{i}=-\infty,\qquad\text{almost surely.}$. (4.31)
4.4. Proving Theorems 4.2, 4.4, and 4.5
In this sub-section, we first prove part $(a)$ of Theorem 4.2, then part $(b)$ of Theorem 4.2 and Theorem 4.4, and we conclude by proving Theorem 4.5.
@@
$\text{Win}(u,v):=\{\exists T\geq 0:\forall t\geq T\ \deg_{t}(u)\geq\deg_{t}(v)\}.$ (4.32)
@@
Assume Assumption 4.1 is satisfied. Then,
@@
    If $\sum_{i=1}^{\infty}X(i)=\infty$ almost surely, then $\lim_{t\to\infty}\max_{v\in\mathrm{BP}(t)}\deg_{t}(v)=\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely.
@@
    If $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ converges almost surely, then for all $u,v\in\mathcal{U}_{\infty}$,
@@
    If $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely, there exists a function $\phi:\mathbb{N}\to\mathbb{N}$ such that $\phi(j)>j$ for all $j\in\mathbb{N}$ such that for any $u\in\mathcal{U}_{\infty}$ and $j\in\mathbb{N}$,
    $\mathbb{P}\!\left(\exists k\in\{j+1,\ldots,\phi(j)\}:\sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji)\right)\geq\frac{1}{2}.$ (4.34)
@@
$(a)$ Let $V$ be the first individual born in $(\mathrm{BP}(t))_{t\geq 0}$ such that $D^{(V)}=\infty$ (taking an arbitrary one in case multiple individuals $u_{1},\ldots,u_{M}$ with $D_{u_{1}}=\cdots=D_{u_{M}}=\infty$ are born simultaneously, for some $M\geq 2$). Conditionally on survival such an individual $V$ exists almost surely, since $\mathbb{P}\!\left(D=\infty\right)>0$. Furthermore, the birth-time of $V$ is finite almost surely. As a result, for all $t\geq\mathcal{B}(V)$,
@@
and the right-hand side tends to infinity with $t$ since $D^{(V)}=\infty$ and as $\sum_{i=1}^{\infty}X(Vi)=\infty$ almost surely (since $X(Vi)\overset{\mathrm{d}}{=}X(i)$ for all $i\in\mathbb{N}$), which yields the desired result.
$(b)$ The event $\text{Win}(u,v)\cup\text{Win}(v,u)$ is trivially satisfied if at most one of the two individuals $u$ and $v$ is born. Given that both vertices are born, again the event is trivially satisfied if at most one individual produces an infinite offspring. As a result, we define
$\mathcal{E}(u,v):=\{\mathcal{B}(u),\mathcal{B}(v)<\infty\}\cap\{D^{(u)},D_{v}=\infty\},$ (4.36)
@@
Recall the assumption that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ converges almost surely. Then, for $u,v\in\mathcal{U}_{\infty}$, on the event $\mathcal{E}(u,v)$ we have
$\left|\mathcal{B}(u)-\mathcal{B}(v)+\sum_{i=1}^{\infty}(X(ui)-X(vi))\right|<\infty\quad\text{almost surely.}$. (4.39)
Assume, without loss of generality, that $|u|\leq|v|$. Let $N$ be chosen such that $\mathcal{B}(v)$ is independent of $(X(un))_{n\geq N}$. For example, if $u$ is a non-parent ancestor of $v$, we may choose the value $N$ such that $u(N-1)$ is the ancestor of $v$, whilst if $v=uj$ for $j\in\mathbb{N}$, we can choose $N>j$. Then, the sequences $(X(un))_{n\geq N}$ and $(X(vn))_{n\geq N}$ are independent. We now have two cases:
(I) Suppose that the sequence $(S_{i})_{i\geq N}$ defined by $S_{i}:=X(ui)-X(vi)$ does not satisfy (4.29). Then, by Lemma 4.9, the random variable
$\sum_{j=N}^{\infty}(X(uj)-X(vj))$ (4.40)
contains no atom on $\mathbb{R}$. Hence, as the summands are independent
$\mathbb{P}\bigg{(}\mathcal{E}(u,v)\cap\bigg{\{}\mathcal{B}(u)-\mathcal{B}(v)+\sum_{j=1}^{\infty}(X(uj)-X(vj))=0\bigg{\}}\bigg{)}=0.$ (4.41)
Therefore, on the event $\mathcal{E}(u,v)$ there exists $K_{0}\in\mathbb{N}$ such that either
$\mathcal{B}(u)-\mathcal{B}(v)+\sum_{j=1}^{k}(X(uj)-X(vj))>0\qquad\text{for all $k\geq K_{0}$,}$ (4.42)
@@
$\mathcal{B}(u)-\mathcal{B}(v)+\sum_{j=1}^{k}(X(uj)-X(vj))<0\qquad\text{for all $k\geq K_{0}$,}$ (4.43)
is satisfied $\mathbb{P}_{\mathcal{S}}$-almost surely. In other words, we either have $\mathcal{B}(uk)>\mathcal{B}(vk)$ for all $k\geq K_{0}$, or we have $\mathcal{B}(uk)<\mathcal{B}(vk)$ for all $k\geq K_{0}$ $\mathbb{P}_{\mathcal{S}}$-almost surely on the event $\mathcal{E}(u,v)$. This in turn implies that for all $k$ sufficiently large $u$ has offspring $k$ before $v$ has offspring $k$ or vice versa. Thus, we have that (4.38) is satisfied.
(II) Suppose that the sequence $(S_{i})_{i\geq N}$ defined by $S_{i}:=X(ui)-X(vi)$ does satisfy (4.29). Since $S_{i}$ has a symmetric distribution (as $X(ui)\overset{\mathrm{d}}{=}X(vi)$), any collection $(c_{i})_{i\in\mathbb{N}}$ in Lemma 4.9 can only contain finitely many non-zero entries (see Remark 4.10). As a result, this implies that $\sum_{i=1}^{\infty}\mathbb{P}\!\left(S_{i}\neq 0\right)=\sum_{i=1}^{\infty}\mathbb{P}(X(ui)\neq X(vi))<\infty$. By the Borel-Cantelli lemma we obtain that $X(ui)=X(vi)$ for all but finitely many $i\in\mathbb{N}$ almost surely. Thus, for some $K_{0}\in\mathbb{N}$, it is either the case that (4.42) or (4.43) is satisfied (but not both), or for all $k\geq K_{0}$,
$\mathcal{B}(u)-\mathcal{B}(v)+\sum_{i=1}^{k}(X(ui)-X(vi))=0.$ (4.44)
This implies that $\mathcal{B}(uk)\geq\mathcal{B}(vk)$ for all $k\geq K_{0}$ or $\mathcal{B}(u_{k})\leq\mathcal{B}(v_{k})$ for all $k\geq K_{0}$. Again, it follows that $\text{Win}(u,v)$ or $\text{Win}(v,u)$ occurs on the event $\mathcal{E}(u,v)$, so that (4.38) holds, which concludes the proof of part $(b)$.
$(c)$ Since $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely, Lemma 4.11 implies that for any $j\in\mathbb{N}$ we have
$\limsup_{n\to\infty}\sum_{i=j+1}^{n}(X(ui)-X(uji))=\infty\quad\text{almost surely}.$ (4.45)
Since also $\sum_{i=1}^{j}X(uji)<\infty$ almost surely by Assumption 4.1, we thus obtain that
$\mathbb{P}\!\left(\exists k>j:\sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji)\right)=1.$ (4.46)
@@
$\lim_{n\to\infty}\mathbb{P}\!\left(\exists k\in\{j+1,\ldots,n\}:\sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji)\right)=1.$ (4.47)
@@
$\phi(j):=\inf\bigg{\{}n>j:\mathbb{P}\bigg{(}\exists k\in\{j+1,\ldots,n\}:\sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji)\bigg{)}\geq\frac{1}{2}\bigg{\}},$ (4.48)
@@
We are now ready to prove part $(a)$ of Theorem 4.2, i.e. that almost surely no persistent $m$-hub emerges $\mathbb{P}_{\mathcal{S}}$-almost surely when the series $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely.
Proof of Theorem 4.2 $(a)$.
@@
where the last step follows from the fact that the event $\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}$ implies survival. We then use that, upon the event $\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}$, $u$ can be a persistent $m$-hub only when for some $T\geq\mathcal{B}(u)$ it has a larger offspring than all but at most $m-1$ many of its children for all time $t\geq T$. That is,
[…]
Here, the second step follows from the fact that any individual almost surely takes an infinite amount of time to produce an infinite offspring by assuming that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely (see Remark 4.3). We thus obtain
$\mathbb{P}_{\mathcal{S}}{}$ $(u\text{ is a persistent hub})$ (4.51)
[…]
We now show that the second probability on the right-hand side equals zero, by showing that
($\mathbb{P}\big{(}\{\exists j\in\mathbb{N}\setminus C\,\exists k\geq K:\mathcal{B}(uk)>\mathcal{B}(ujk)\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big{)}=\mathbb{P}\big{(}\mathcal{B}(u)<\infty,D^{(u)}=\infty\big{)}.$ (4.52)
On the event $\{\mathcal{B}(u)<\infty\}$, we can write
$\mathcal{B}(uk)>\mathcal{B}(ujk)\ \ \Leftrightarrow\ \ \sum_{i=1}^{k}X(ui)>\sum_{i=1}^{j}X(ui)+\sum_{i=1}^{k}X(uji)\ \ \Leftrightarrow\ \ \sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji).$ (4.53)
We now recall the function $\phi$ in part $(c)$ of Lemma 4.12. We let $(j_{\ell})_{\ell\in\mathbb{N}}=j_{\ell}(C,K)$ be the subsequence $j_{1}=1+\max\{\max C,K\}$ and $j_{\ell+1}:=\phi(j_{\ell})$ for all $\ell\in\mathbb{N}$. As $\phi(j)>j$ for all $j\in\mathbb{N}$, it follows that the subsequence $(j_{\ell})_{\ell\in\mathbb{N}}$ is strictly increasing and thus $j_{\ell}>\max C$ and $j_{\ell}>K$ for all $\ell\in\mathbb{N}$. Then,
$\mathbb{P}{}$ ($\big{(}\{\exists j\in\mathbb{N}\setminus C\,\exists k\geq K:\mathcal{B}(uk)>\mathcal{B}(ujk)\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big{)}$ (4.54)
=($=\mathbb{P}\bigg{(}\bigg{\{}\exists j\in\mathbb{N}\setminus C\,\exists k\geq K:\sum_{i=j+1}^{k}X(ui)>\sum_{i=1}^{k}X(uji)\bigg{\}}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\bigg{)}$
($\geq\mathbb{P}\bigg{(}\bigg{\{}\exists\ell\in\mathbb{N}\,\exists k\in\{j_{\ell}+1,\ldots,\phi(j_{\ell})\}:\!\!\!\sum_{i=j_{\ell}+1}^{k}\!\!\!X(ui)>\sum_{i=1}^{k}X(uj_{\ell}i)\bigg{\}}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\bigg{)}.$
For ease of writing, we define
$\mathcal{E}(u,\ell):=\bigg{\{}\exists k\in\{j_{\ell}+1,\ldots,\phi(j_{\ell})\}:\sum_{i=j_{\ell}+1}^{k}X(ui)>\sum_{i=1}^{k}X(uj_{\ell}i)\bigg{\}}\qquad\text{for }\ell\in\mathbb{N}.$ (4.55)

### v2

@@ 4.2. Preliminary results
We first collect a few preliminary results that are used throughout the proof of Theorem 4.1.
@@
$\mathbb{P}\bigg(\exists j\in\mathbb{N}_{0}\colon Y+\sum_{i=1}^{k+j}X(i)\leq\sum_{i=k}^{k+j}X(i)^{\prime}\bigg)\leq\bigg(\prod_{i=k}^{\infty}\mathbb{E}\Big[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\Big]\bigg)\mathbb{E}\bigg[\exp\bigg(-\lambda\bigg(Y+\sum_{i=1}^{k-1}X(i)\bigg)\bigg)\bigg].$ (4.9)
In short, the proof of Lemma 4.3 is based on the fact that the random variable $\exp(\lambda\sum_{i=k}^{k+j}(X(i)-X(i)^{\prime}))$ is a sub-martingale combined with Chernoff’s inequality.
Lemma 4.4 (Proposition $1.15$ in [24]).
Suppose that $(S_{i})_{j\in\mathbb{N}}$ is a sequence of mutually independent, symmetric random variables. If $\sum_{i=1}^{\infty}S_{i}$ diverges almost surely, then
$\limsup_{n\to\infty}\sum_{i=1}^{n}S_{i}=\infty\qquad\text{and}\qquad\liminf_{n\to\infty}\sum_{i=1}^{n}S_{i}=-\infty,\qquad\text{almost surely.}$ (4.10)
The proof of Lemma 4.4 is based on the fact that $\sum_{i=1}^{n}S_{i}$ is a martingale, combined with the fact that divergent martingales are not bounded almost surely.
Lemma 4.5.
Let $(E_{i})_{i\in\mathbb{N}_{0}}$ and $(E_{i}^{\prime})_{i\in\mathbb{N}_{0}}$ be independent exponential random variables, where $E_{i}$ and $E_{i}^{\prime}$ have rate $\lambda_{i}>0$ for each $i\in\mathbb{N}_{0}$. Then,
$\sum_{i=0}^{\infty}(E_{i}-E_{i}^{\prime})\text{ converges almost surely}\quad\Leftrightarrow\quad\sum_{i=0}^{\infty}\frac{1}{\lambda_{i}^{2}}<\infty.$ (4.11)
@@
$(i)$ If we set $X(i)=E_{i}$ and $X(i)^{\prime}=E_{i}^{\prime}$ in Lemma 4.3, it also follows that for any $\lambda>0$ there exist $k\in\mathbb{N}$ such that (4.8) is satisfied when
$\sum_{i=0}^{\infty}\frac{1}{\lambda_{i}^{2}}<\infty.$ (4.12)
Indeed, as the sum is finite it follows that $\lambda_{i}\to\infty$ with $i$, so that for any $\lambda>0$ we can take $k$ large enough such that $\lambda_{i}>\lambda$ for all $i\geq k$, so that
$\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]=\frac{\lambda_{i}^{2}}{(\lambda_{i}-\lambda)(\lambda_{i}+\lambda)}=\frac{1}{1-\lambda^{2}/\lambda_{i}^{2}}.$ (4.13)
The product in (4.8) hence is finite when the sum in (4.12) is finite.
[…]
Proof.
To prove the convergence of the series of exponential random variables, we check the conditions of Kolmogorov’s three series theorem (see e.g. [19, Theorem $2.5.4$]). This theorem states that a random series $\sum_{i=0}^{\infty}X_{i}$ of independent random variables $(X_{i})_{i\in\mathbb{N}_{0}}$ converges almost surely if the following conditions holds for some $A>0$, and only if the following conditions hold for any $A>0$:
[…]
We notice that for symmetric random variables, the second condition is always satisfied. It thus suffices to focus on conditions $(i)$ and $(iii)$, for $X_{i}=E_{i}-E_{i}^{\prime}$.
Let us first prove the ‘if’ direction. We suppose that $\sum_{i=0}^{\infty}\lambda_{i}^{-2}<\infty$. Then,
$\sum_{i=0}^{\infty}{\rm Var}(X_{i}\mathbbm{1}_{\{|X_{i}|\leq A\}})=\sum_{i=0}^{\infty}\mathbb{E}\left[X_{i}^{2}\mathbbm{1}_{\{|X_{i}|\leq A\}}\right]\leq\sum_{i=0}^{\infty}\mathbb{E}\left[X_{i}^{2}\right]=\sum_{i=0}^{\infty}2{\rm Var}(E_{i})=\sum_{i=0}^{\infty}\frac{2}{\lambda_{i}^{2}},$ (4.15)
which is finite by assumption. Condition $(iii)$ is hence satisfied. Also,
$\sum_{i=0}^{\infty}\mathbb{P}\!\left(|X_{i}|\geq A\right)\leq\sum_{i=0}^{\infty}\mathbb{P}\!\left(E_{i}+E_{i}^{\prime}\geq A\right)\leq 2\sum_{i=0}^{\infty}\mathbb{P}\!\left(E_{i}\geq\frac{A}{2}\right)=2\sum_{i=0}^{\infty}\mathrm{e}^{-\lambda_{i}A/2}$ (4.16)
By assuming that $\sum_{i=0}^{\infty}\lambda_{i}^{-2}<\infty$, it follows that $\lambda_{i}\to\infty$ with $i$. Hence, there exists $I\in\mathbb{N}$ such that $\mathrm{e}^{-\lambda_{i}A/2}\leq 1/\lambda_{i}^{2}$, from which Condition $(i)$ follows.
For the ‘only if’ direction, we argue by contradiction. We assume that there exists $A>0$ such that either Condition $(i)$ or $(iii)$ does not hold, and additionally assume that $\sum_{i=0}^{\infty}\lambda_{i}^{-2}<\infty$. We show that this leads to a contradiction.
Suppose first that Condition $(i)$ is not satisfied for some $A>0$. By the bound in (4.16) and the assumption that $\sum_{i=0}^{\infty}\lambda_{i}^{-2}<\infty$, this directly leads to a contradiction. Hence, now assume that Condition $(i)$ is satisfied for all $A>0$, but that Condition $(iii)$ is not satisfied for some $A>0$. Then, by the second Borel-Cantelli lemma, $|X_{i}|\geq A$ only finitely often almost surely, so that
$\sum_{i=0}^{\infty}{\rm Var}(X_{i}\mathbbm{1}_{\{|X_{i}|\leq A\}})=\infty\quad\Leftrightarrow\quad\sum_{i=0}^{\infty}{\rm Var}(X_{i})=\infty.$ (4.17)
However, the equality on the right-hand side contradicts the assumption that $\sum_{i=0}^{\infty}\lambda_{i}^{-2}<\infty$, since ${\rm Var}(X_{i})=\mathbb{E}\left[X_{i}^{2}\right]=2{\rm Var}{E_{i}}=2\lambda_{i}^{-2}$. ∎
4.3. Proving Theorem 4.1(a)
@@
$\text{Win}(u,v)\coloneq\{\exists T\geq 0:\forall t\geq T\ \deg_{t}(u)\geq\deg_{t}(v)\}.$ (4.18)
@@
Recall the CTPAVD branching process $(\mathrm{BP}(t))_{t\geq 0}$ defined in (3.4), with offspring, inter-birth times, and lifetimes as defined in (the paragraph before) (3.7). Assume that Assumption (N-E) is satisfied and Assumption (F-D) is not satisfied. The following statements are true:
@@
    $\lim_{t\to\infty}\max_{v\in\mathrm{BP}(t)}\deg_{t}(v)=\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely.
@@
    If Assumption (C-V) is satisfied, then for all $u,v\in\mathcal{U}_{\infty}$,
@@
    If Assumption (C-V) is not satisfied, there exists a function $\phi:\mathbb{N}\to\mathbb{N}$ such that $\phi(j)>j$ for all $j\in\mathbb{N}$ such that for any $u\in\mathcal{U}_{\infty}$ and $j\in\mathbb{N}$,
    $\mathbb{P}\!\left(\exists k\in\{j+1,\ldots,\phi(j)\}:\sum_{i=j}^{k-1}E^{(u)}_{i}>\sum_{i=0}^{k-1}E^{(uj)}_{i}\right)\geq\frac{1}{2}.$ (4.20)
@@
$(a)$ As Assumption (F-D) is not satisfied, it follows from [23, Lemma 5.3 and Corollary 6.4] that offspring random variable $D$ is such that $\mathbb{P}\!\left(D=\infty\right)>0$. We can hence define $V$ be the first individual born in $(\mathrm{BP}(t))_{t\geq 0}$ such that $D^{(V)}=\infty$. Conditionally on survival such an individual $V$ exists almost surely. Furthermore, the birth-time of $V$ is finite almost surely. As a result, for all $t\geq\mathcal{B}(V)$,
@@
and the right-hand side tends to infinity with $t$ since $D^{(V)}=\infty$.
$(b)$ The event $\text{Win}(u,v)\cup\text{Win}(v,u)$ is trivially satisfied if at most one of the two individuals $u$ and $v$ is born. Given that both vertices are born, again the event is trivially satisfied if at most one individual produces infinite offspring. As a result, we define
$\mathcal{E}(u,v)\coloneq\{\mathcal{B}(u),\mathcal{B}(v)<\infty\}\cap\{D^{(u)},D_{v}=\infty\},$ (4.22)
@@
Lemma 4.5 tells us that Assumption (C-V) being satisfied is equivalent to the infinite series $\sum_{i=0}^{\infty}(E_{i}-E_{i}^{\prime})$ converging almost surely. Then, for $u,v\in\mathcal{U}_{\infty}$, on the event $\mathcal{E}(u,v)$ we have
$\left|\mathcal{B}(u)-\mathcal{B}(v)+\sum_{i=0}^{\infty}(E^{(u)}_{i}-E^{(v)}_{i})\right|<\infty\quad\text{almost surely,}$ (4.25)
since $E^{(u)}_{i}\overset{\mathrm{d}}{=}E_{i}$ and $E^{(v)}_{i}\overset{\mathrm{d}}{=}E_{i}^{\prime}$. Assume, without loss of generality, that $|u|\leq|v|$. Let $N$ be chosen such that $\mathcal{B}(v)$ is independent of $(E^{(u)}_{n})_{n\geq N}$. For example, if $u$ is a non-parent ancestor of $v$, we may choose the value $N$ such that u(N1)u(N-1) is the ancestor of $v$, whilst if $v=uj$ for $j\in\mathbb{N}$, we can choose $N>j$. Then, the sequences $(E^{(u)}_{n})_{n\geq N}$ and $(E^{(v)}_{n})_{n\geq N}$ are independent. Therefore, on the event $\mathcal{E}(u,v)$ there exists $K_{0}\in\mathbb{N}$ such that either
$\mathcal{B}(u)-\mathcal{B}(v)+\sum_{j=0}^{k-1}(E^{(u)}_{j}-E^{(v)}_{j})>0\qquad\text{for all $k\geq K_{0}$,}$ (4.26)
@@
$\mathcal{B}(u)-\mathcal{B}(v)+\sum_{j=0}^{k-1}(E^{(u)}_{j}-E^{(v)}_{j})<0\qquad\text{for all $k\geq K_{0}$,}$ (4.27)
is satisfied $\mathbb{P}_{\mathcal{S}}$-almost surely. In other words, we either have $\mathcal{B}(uk)>\mathcal{B}(vk)$ for all $k\geq K_{0}$, or we have $\mathcal{B}(uk)<\mathcal{B}(vk)$ for all $k\geq K_{0}$ $\mathbb{P}_{\mathcal{S}}$-almost surely on the event $\mathcal{E}(u,v)$. This in turn implies that for all $k$ sufficiently large $u$ has offspring $k$ before $v$ has offspring $k$ or vice versa. Thus, we have that (4.24) is satisfied, which concludes the proof of part $(b)$.
$(c)$ We again use Lemma 4.5 to conclude that Assumption (C-V) not being satisfied is equivalent to the infinite series $\sum_{i=0}^{\infty}(E_{i}-E_{i}^{\prime})$ diverging almost surely, and use again that $E^{(u)}_{i}\overset{\mathrm{d}}{=}E_{i}$ and $E^{(uj)}_{i}\overset{\mathrm{d}}{=}E_{i}^{\prime}$ for any $j\in\mathbb{N}$ and any $u\in\mathcal{U}_{\infty}$. We then use Lemma 4.4 to obtain that for any $j\in\mathbb{N}$ we have
$\limsup_{n\to\infty}\sum_{i=j}^{n}(E^{(u)}_{i}-E^{(uj)}_{i})=\infty\quad\text{almost surely}.$ (4.28)
Since also $\sum_{i=0}^{j-1}E^{(u)}_{i}$$\infty$ almost surely, we thus obtain that
$\mathbb{P}\!\left(\exists k>j:\sum_{i=j}^{k}E^{(u)}_{i}>\sum_{i=0}^{k}E^{(uj)}_{i}\right)=1.$ (4.29)
@@
$\lim_{n\to\infty}\mathbb{P}\!\left(\exists k\in\{j+1,\ldots,n\}:\sum_{i=j}^{k-1}E^{(u)}_{i}>\sum_{i=0}^{k-1}E^{(uj)}_{i}\right)=1.$ (4.30)
@@
$\phi(j)\coloneq\inf\bigg\{n>j:\mathbb{P}\bigg(\exists k\in\{j+1,\ldots,n\}:\sum_{i=j}^{k-1}E^{(u)}_{i}>\sum_{i=0}^{k-1}E^{(uj)}_{i}\bigg)\geq\frac{1}{2}\bigg\},$ (4.31)
@@
We are now ready to prove part $(a)$ of Theorem 4.1, i.e. that almost surely no persistent $m$-hub emerges $\mathbb{P}_{\mathcal{S}}$-almost surely when Assumption (C-V) is not satisfied (i.e. when the series $\sum_{i=0}^{\infty}(E_{i}-E_{i}^{\prime})$ diverges almost surely.
Proof of Theorem 4.1(a)(a).
@@
$\mathbb{P}_{\mathcal{S}}\left(u\text{ is a persistent $m$-hub}\right)$ $=\mathbb{P}_{\mathcal{S}}\big(\{u\text{ is a persistent $m$-hub}\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big)$ (4.32)
where the last step follows from the fact that the event $\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}$ implies survival. We then use that, upon the event $\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}$, $u$ can be a persistent $m$-hub only when for some $T\geq\mathcal{B}(u)$ it has a larger offspring than all but at most m1m-1 many of its children for all time $t\geq T$. That is,
$\{u\text{ is a persistent $m$-hub}\}$ $\subseteq\bigcup_{T\geq\mathcal{B}(u)}\bigcup_{\begin{subarray}{c}C\subset\mathbb{N}\\ |C|<m\end{subarray}}\{\forall j\in\mathbb{N}\setminus C\ \forall t\geq T:\deg_{t}(u)\geq\deg_{t}(uj)\}$ (4.33)
Here, the second step follows from the fact that any individual almost surely takes an infinite amount of time to produce infinite offspring by Assumption (N-E). We thus obtain
$\mathbb{P}_{\mathcal{S}}{}$ $(u\text{ is a persistent hub})$ (4.34)
$\leq{}$ $\frac{1}{\mathbb{P}\!\left(\mathcal{S}\right)}\sum_{K=1}^{\infty}\sum_{\begin{subarray}{c}C\subset\mathbb{N}\\ |C|<m\end{subarray}}\mathbb{P}\big(\{\forall j\in\mathbb{N}\setminus C\ \forall k\geq K:\mathcal{B}(uk)\leq\mathcal{B}(ujk)\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big).$
We now show that each probability on the right-hand side equals zero, by showing that
$\mathbb{P}\big(\{\exists j\in\mathbb{N}\setminus C\,\exists k\geq K:\mathcal{B}(uk)>\mathcal{B}(ujk)\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big)=\mathbb{P}\big(\mathcal{B}(u)<\infty,D^{(u)}=\infty\big).$ (4.35)
On the event $\{\mathcal{B}(u)<\infty\}$, we can write
$\mathcal{B}(uk)>\mathcal{B}(ujk)\ \ \Leftrightarrow\ \ \sum_{i=0}^{k-1}E^{(u)}_{i}>\sum_{i=0}^{j-1}E^{(u)}_{i}+\sum_{i=0}^{k-1}E^{(uj)}_{i}\ \ \Leftrightarrow\ \ \sum_{i=j}^{k-1}E^{(u)}_{i}>\sum_{i=0}^{k-1}E^{(uj)}_{i}$. (4.36)
We now recall the function $\phi$ in part $(c)$ of Lemma 4.7. We let $(j_{\ell})_{\ell\in\mathbb{N}}=j_{\ell}(C,K)$ be the subsequence $j_{1}=1+\max\{\max C,K\}$ and $j_{\ell+1}\coloneq\phi(j_{\ell})$ for all $\ell\in\mathbb{N}$. As $\phi(j)>j$ for all $j\in\mathbb{N}$, it follows that the subsequence $(j_{\ell})_{\ell\in\mathbb{N}}$ is strictly increasing and thus $j_{\ell}>\max C$ and $j_{\ell}>K$ for all $\ell\in\mathbb{N}$. Then,
$\mathbb{P}{}$ $\big(\{\exists j\in\mathbb{N}\setminus C\,\exists k\geq K:\mathcal{B}(uk)>\mathcal{B}(ujk)\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big)$ (4.37)
$=\mathbb{P}\bigg(\bigg\{\exists j\in\mathbb{N}\setminus C\,\exists k\geq K:\sum_{i=j}^{k-1}E^{(u)}_{i}>\sum_{i=0}^{k-1}E^{(uj)}_{i}\bigg\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\bigg)$
$\geq\mathbb{P}\bigg(\bigg\{\exists\ell\in\mathbb{N}\,\exists k\in\{j_{\ell}+1,\ldots,\phi(j_{\ell})\}\colon\sum_{i=j_{\ell}}^{k-1}E^{(u)}_{i}>\sum_{i=0}^{k-1}E^{(uj_{\ell})}_{i}\bigg\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\bigg).$
For ease of writing, we define